# Beyond Fidelity: Explaining Vulnerability Localization of Learning-based Detectors

### 05 Jan 2024 v1

@@ 8. Conclusion
In this study, we systematically evaluate the challenges of DL-based detectors in locating vulnerabilities through explanation methods. We find that these methods have limited effectiveness in solving the problem. Two factors contribute to this issue. Firstly, the explainers may not accurately capture the vulnerability-related features. Secondly, the DL-based detectors may learn artifacts unrelated to the vulnerability, as retraining models on a balanced dataset can indeed improve model-related explainers’ performance regarding FLC. This could be attributed to the occurrence of some vulnerability-unrelated distinctions between vulnerable and normal samples.

### 21 Feb 2024 v2

@@ 8. Conclusion
In this study, we systematically evaluate the challenges in locating vulnerabilities with DL-based detectors and explanation approaches. Firstly, we evaluate the effectiveness of the explanation approaches using two metrics: fidelity and LC. We find that the results of fidelity and LC show little correlation. Meanwhile, these explanation methods show limited effectiveness in locating vulnerable lines. Secondly, subsequent experiments with DL-based detectors suggest that they might learn artifacts unrelated to the vulnerabilities, as retraining models on a balanced dataset can indeed improve model-related explainers’ performance regarding FLC. This phenomenon may stem from vulnerability-unrelated distinctions between vulnerable and normal samples, thereby diminishing the efficacy of explanation approaches in identifying vulnerable lines. Furthermore, even when trained on a balanced dataset, models remain susceptible to vulnerability-unrelated perturbations, undermining the reliability of fidelity calculations or explanation approaches based on perturbation. Based on our findings, employing binary classification models directly for vulnerability detection may not be effective. We suggest that future work endeavors decompose vulnerability detection into multiple subtasks and utilize deep learning techniques to improve efficiency.
Acknowledgements.
This work was partially supported by the National Natural Science Foundation of China (No. 62141208), the National Natural Science Foundation of China (grant No.62072046), the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), HUST CSE-HongXin Joint Institute for Cyber Security, and HUST CSE-FiberHome Joint Institute for Cyber Security, and the Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the respective funding agencies.